# Relaxation of weakly collisional plasma: continuous spectra, discrete eigenmodes, and the decay of echoes

### v2

@@ Abstract
The relaxation of a weakly collisional plasma, which is of fundamental interest to laboratory and fusion plasmas as well as astrophysical plasmas, can be described by the self-consistent Boltzmann-Poisson equations with the Lenard-Bernstein collision operator. We perform a perturbative (linear and second order) analysis of the Boltzmann-Poisson equations, and obtain exact analytic solutions which enable definitive resolutions to some long-standing controversies regarding the impact of weak collisions on the continuous spectra, the discrete Landau eigenmodes, and the decay of plasma echoes. We retain both damping and diffusion terms in the collision operator throughout our treatment. We find that the linear response is a temporal convolution of two types of contribution: a continuum that depends on the continuous velocities of particles (crucial for the plasma echo), and another, consisting of discrete modes that are coherent modes of oscillation of the entire system. The discrete modes are exponentially damped over time due to collective effects or wave-particle interactions (Landau damping), as well as collisional dissipation. The continuum is also damped by collisions, but somewhat differently than the discrete modes. Up to a collision time, which is the inverse of the collision frequency $\nu_{\rm c}$, the continuum decay is driven by the diffusion of particle velocities and is cubic exponential, occurring over a timescale $\sim\nu^{-1/3}_{\rm c}$. After a collision time, however, the continuum decay is driven by the collisional damping of particle velocities and diffusion of their positions, and occurs exponentially over a timescale $\sim\nu_{\rm c}$. This slow exponential decay causes perturbations to damp the most on scales comparable to the mean free path, but very slowly on larger scales. This establishes the local thermal equilibrium (LTE), which is the essence of the fluid limit. The long-term decay of the linear response is driven by the discrete modes on scales smaller than the mean free path, but, on larger scales, is governed by a combination of the slowly decaying continuum and the least damped discrete mode. This slow exponential decay implies that the echo, which results from the interference of the continuum response to two subsequent pulses, is detectable even on scales comparable to the mean free path, as long as the second pulse is introduced within a few phase-mixing timescales after the first.
@@ I Introduction
Understanding the modal response of a weakly collisional plasma is critical to the development of a statistical description of the system (be it in equilibrium or non-equilibrium). Recently, there has been a resurgence of interest in the long-time behavior of weakly collisional kinetic plasmas. A rigorous linear theory of weakly collisional plasmas [9, 10, 11], which builds on the seminal work of [12], demonstrated that a plasma governed by the Lenard-Bernstein (LB) collision operator (a special case of the Fokker-Planck operator) oscillates in the singular limit of zero collisions as a complete set of discrete normal modes at long times. A striking finding of these studies was that the Case-van Kampen modes, which are known to be a complete set of normal modes for a purely collisionless plasma and constitute a continuous spectrum, are eliminated when weak collisions are introduced into the system. The new eigenmodes of a weakly collisional plasma constitute a discrete spectrum, and now include the Landau solutions. As is well known, the Landau solutions are not eigenmodes of the collisionless Vlasov-Poisson system and are instead remnants of the initial perturbations in the time-asymptotic limit. However, in the presence of weak collisions, these become eigenmodes (referred to hereafter as Landau modes [^1] In the collisionless limit, on scales smaller than the Debye length, these are Landau damped, while, on scales larger than the Debye length, these are nothing but the weakly damped electron Langmuir waves and ion acoustic waves.). For a stable plasma, these eigenmodes undergo a simple exponential decay with time. On the other hand, [13] showed in a classic paper that the linear response to perturbations decays cubic exponentially due to collisions (at least at short time). Although they do not make it explicit in their paper, their analysis only applies to the continuum part of the plasma response, as we clarify in section IV. Their findings imply that the plasma echo, which arises non-linearly from the continuum response, would rapidly decay due to electron-ion collisions. The [13] prediction for the temporal decay of the continuum response is, however, much more rapid than the exponential decay of the discrete eigenmodes, identified by [9, 10]. [14] revisited the [13] result from a boundary-layer analysis of the problem, reaffirming the results of [13].
@@
It thus appears that the relaxation of weakly collisional plasmas has produced somewhat conflicting results: the super-exponential (cubic exponential) decay of the continuum response and the plasma echo according to [13] and the exponential decay of the discrete eigenmodes according to [12, 9, 10, 15, 11]. This naturally raises the following questions. Which of these results is correct? And, if both are correct, are these results simply different limits of the same response? This paper addresses these very questions. We show that the exact solution to the problem of complete perturbative (linear and second order) relaxation of a weakly collisional plasma indeed yields the above results as different asymptotic limits in time. We show that including the damping (also known as dynamical friction/drag [16]) term in the collision operator modifies the [13] prediction of cubic exponential decay of the continuum response to an exponential decay at late times and large spatial scales. We find that the [13] prediction only holds for short times, i.e., $t<1/\nu_{\rm c}$, and scales smaller than the mean free path, i.e., $\lambda<\lambda_{\rm c}=\sigma/\nu_{\rm c}$ or $\nu_{\rm c}<k\sigma$, where $\nu_{\rm c}$ is the collision frequency, $\sigma$ is the equilibrium velocity dispersion, $\lambda_{\rm c}=\sigma/\nu_{\rm c}$ is the mean free path, $k$ is the wavenumber of the response and $\lambda\sim 1/k$ is the wavelength. At long times ($t>1/\nu_{\rm c}$) and on large scales ($\lambda>\lambda_{\rm c}$ or $\nu_{\rm c}>k\sigma$), the response damps exponentially and much slower, at a rate similar to that predicted by [9, 10]. The damping is most pronounced for $\nu_{\rm c}\sim k\sigma$ or $\lambda\sim\lambda_{\rm c}$. The slow exponential decay outside the mean free path implies that the distribution function (DF) approaches the equilibrium Maxwellian form rapidly only within the mean free path; this is consistent with the assumption of the establishment of a local thermodynamic equilibrium (LTE) in the fluid approximation (large $\nu_{\rm c}$). Hence, we recover the fluid limit from a rigorous theory of weakly collisional plasmas.
@@ III Solving the linearized equations
where $\nu_{\rm c}$ is the collision frequency and $\sigma$ is the velocity dispersion of the system when it is in equilibrium. The Boltzmann equation with the LB operator is also known as the mean field Kramers equation [see 23, for a detailed discussion of the initial value problem for this equation in the context of weakly collisional plasmas and self-gravitating systems]. The fundamental assumption behind this form of the collision operator is that the collisions are sufficiently short-range and random. The LB operator conserves the particle number but not the linear momentum and kinetic energy during collisions. This shortcoming is well-known in the plasma literature [12, 9, 10, 11]. Despite this shortcoming, we expect the scaling behaviours of different aspects of plasma relaxation to be qualitatively robust (see section IV.2 for details).
@@ IV.1.1 Gaussian pulse
Conservation of particle number ensures that the pulse damps out as well as broadens due to collisions, as manifested by both $x$ and $v$ parts of the above response. The spatial broadening of the pulse occurs as $\sim\left(2\sigma^{2}/3\nu_{\rm c}^{2}\right){\left(\nu_{\rm c}t\right)}^{3}$ for $t\lesssim 1/\nu_{\rm c}$, and as $\sim\left(2\sigma^{2}/\nu_{\rm c}^{2}\right)\nu_{\rm c}t$ for $t\gtrsim 1/\nu_{\rm c}$, while the spatial damping occurs as the inverse of this. If the width of the initial pulse is small, i.e., $\sigma_{x,0}\ll\sigma/\nu_{\rm c}$, it damps and broadens in the former way for a long time, i.e., we are in the diffusive regime. The typical damping timescale in this regime is $\sim{\left(3\sigma^{2}_{x,0}/2\sigma^{2}\nu_{\rm c}\right)}^{1/3}$. On the other hand, a well-extended initial pulse with $\sigma_{x,0}\gg\sigma/\nu_{\rm c}$, falls in the damping regime, i.e., it quickly transitions to the latter phase of exponential damping and broadening, which occurs over a relatively long timescale of $\sim\left(\sigma^{2}_{x,0}/2\sigma^{2}\right)\nu_{\rm c}$. Since both timescales increase with $\sigma_{x,0}$ (for fixed $\nu_{\rm c}$ and $\sigma$), we see that small scale perturbations collisionally damp away faster than large scale ones. Moreover, since the former timescale decreases with $\nu_{\rm c}$ (for fixed $\sigma_{x,0}$ and $\sigma$) and the latter increases, perturbations damp slowly in both small and large $\nu_{\rm c}$ limits. The damping and broadening of the pulse due to collisions is visible in Fig. 2, which plots $f_{1}$ (solid lines) as a function of $x$ at different points of time as indicated. Note that the pulse slows down over time due to collisional damping. The three key features of Fokker-Planck collisional dynamics are the damping, broadening and deceleration of a pulse, all three of which are manifest in Fig. 2.
@@ IV.2 Langevin approach: diffusive and friction regimes
where we have used the uncorrelated nature of the stochastic force, i.e., the second of equations (32). Note that the velocity dispersion initially increases linearly with time, i.e., $\sigma^{2}(t)\sim 2{\cal D}\,t$. This is the essence of velocity diffusion or particle random walk in the velocity space, and is caused by the stochastic acceleration, $\xi(t)$. The late-time dynamics, which sets in after $t\sim 1/\nu_{\rm c}$, is governed by the frictional deceleration, $-\nu_{\rm c}{\rm d}x/{\rm d}t$. This friction/damping drives $\sigma^{2}$ towards the equilibrium value, $\lim_{t\to\infty}\sigma^{2}(t)=\sigma^{2}$. Thus, from the above equation, we infer that ${\cal D}=\nu_{\rm c}\sigma^{2}$. This is nothing but the celebrated Einstein relation. ${\cal D}$ is the velocity diffusion coefficient, which is related to the damping coefficient or collision frequency, $\nu_{\rm c}$, by the proportionality factor, $\sigma^{2}$. Therefore, friction and diffusion go hand-in-hand during collisions, and one cannot exist without the other. Consequently, neglecting the friction term in the collision operator [13, 14] is bound to provide an incomplete description of collisional dynamics, which may lead to erroneous conclusions about the long time decay of the response in weakly collisional plasmas.
@@ V Linear response: discrete modes
So far, we have studied the first part of the plasma response, the continuum, which encapsulates the collisionally damped streaming of particles. The second part of the response on the other hand consists of discrete modes [^3] For each $k$, there are countably many modes. that are collective coherent oscillations of the system. It can be computed by taking the inverse Laplace transform of $\tilde{G}_{1\kappa}^{(2)}(\omega)$ in equation (61). The resulting response involves a convolution of the continuum and discrete mode responses. We defer the computation of the combined response to the next section, and only discuss the behaviour of the discrete modes in this section. The discrete modes oscillate at frequencies that follow the dispersion relation,
@@ V.2 Lenard-Bernstein Branch
The collision operator in the Boltzmann equation involves the highest velocity derivative and is therefore a singular perturbation to the collisionless Boltzmann equation. This implies that the discrete mode spectrum for the weakly collisional case is qualitatively different from the collisionless case. Therefore, in addition to the Landau branch modes discussed above, the LB operator introduces an additional set of modes, which we term Lenard-Bernstein or LB modes. We indicate these using white dashed lines in Fig. 3. Unlike the Landau branch modes, the LB modes do not analytically continue to the Landau modes in the collisionless limit. Moreover, the LB modes are damped and non-oscillatory. As also found by [12], the frequencies of these modes are given by
@@
The LB modes play a crucial role in plasma relaxation in the large $\nu_{\rm c}/k\sigma$ limit. The damping timescale of the LB modes scales with $\nu_{\rm c}$ in the same way as that of the Landau branch modes in this limit. Also, in this limit, the damping time of the continuum response is the same as that of the $n=0$ LB mode. Hence, it is the LB modes that dictate the relaxation of plasmas in the limit of frequent collisions, or equivalently on scales larger than the mean free path.
Vlasov simulations have demonstrated some of the phenomenology of wave damping described in this paper. [27] find in their simulation the survival of a non-oscillating and slowly decaying ‘entropy mode’ (also discussed by [21]) in the long-time limit. This bears qualitative resemblance to the LB modes [12] discussed in this paper. The long term, slow decay of the response in their simulation is very similar to that of the slowly decaying continuum and the $n=0$ LB mode that we predict (although non-linear Landau damping effects become important after $\sim$ the bounce time or libration time for the wave-particle resonances). The pronounced damping of the response for $\nu_{\rm c}\sim k\sigma$, as predicted by our analysis, is seen by [28] in their Vlasov simulations, who note that the collision frequency suitable for preventing the “numerical recurrence of initial conditions” (the echo; see section VII for a detailed discussion of how echoes damp) is a function of the wavenumber of the perturbation. We show that this wavenumber is nothing but $\sim\nu_{\rm c}/\sigma$, the inverse of the mean free path. Perturbations tend to damp away cubic exponentially within a mean free path, and undergo slow, exponential damping beyond.

### v3

@@ Abstract
The relaxation of a weakly collisional plasma, which is of fundamental interest to laboratory and astrophysical plasmas, can be described by the self-consistent Boltzmann-Poisson equations with the Lenard-Bernstein collision operator. We perform a perturbative (linear and second order) analysis of the Boltzmann-Poisson equations, and obtain exact analytic solutions which resolve some long-standing controversies regarding the impact of weak collisions on the continuous spectra, the discrete Landau eigenmodes, and the decay of plasma echoes. We retain both damping and diffusion terms in the collision operator throughout our treatment. We find that the linear response is a temporal convolution of two types of contribution: a continuum that depends on the continuous velocities of particles (crucial for the plasma echo), and another, consisting of discrete modes that are coherent modes of oscillation of the entire system. The discrete modes are exponentially damped over time due to collective effects or wave-particle interactions (Landau damping), as well as collisional dissipation. The continuum is also damped by collisions, but somewhat differently than the discrete modes. Up to a collision time, which is the inverse of the collision frequency $\nu_{\rm c}$, the continuum decay is driven by the diffusion of particle velocities and is cubic exponential, occurring over a timescale $\sim\nu^{-1/3}_{\rm c}$. After a collision time, however, the continuum decay is driven by the collisional damping of particle velocities and diffusion of their positions, and occurs exponentially over a timescale $\sim\nu_{\rm c}$. This slow exponential decay causes perturbations to damp the most on scales comparable to the mean free path, but very slowly on larger scales. This establishes the local thermal equilibrium (LTE), which is the essence of the fluid limit. The long-term decay of the linear response is driven by the discrete modes on scales smaller than the mean free path, but, on larger scales, is governed by a combination of the slowly decaying continuum and the least damped discrete mode. This slow exponential decay implies that the echo, which results from the interference of the continuum response to two subsequent pulses, is detectable even on scales comparable to the mean free path, as long as the second pulse is introduced within a few phase-mixing timescales after the first.
@@ I Introduction
Understanding the modal response of a weakly collisional plasma is critical to the development of a statistical description of the system (be it in equilibrium or non-equilibrium). Recently, there has been a resurgence of interest in the long-time behavior of weakly collisional kinetic plasmas. A rigorous linear theory of weakly collisional plasmas [9, 10, 11], which builds on the seminal work of [12], demonstrated that a plasma governed by the Lenard-Bernstein (LB) collision operator (a special case of the Fokker-Planck operator) oscillates in the singular limit of zero collisions as a complete set of discrete normal modes at long times. A striking finding of these studies was that the Case-van Kampen modes, which are known to be a complete set of normal modes for a purely collisionless plasma and constitute a continuous spectrum, are eliminated when weak collisions are introduced into the system. The new eigenmodes of a weakly collisional plasma constitute a discrete spectrum, and include the Landau solutions. As is well known, the Landau solutions are not eigenmodes of the collisionless Vlasov-Poisson system and are instead remnants of the initial perturbations in the time-asymptotic limit. However, in the presence of weak collisions, these become eigenmodes (referred to hereafter as Landau modes [^1] In the collisionless limit, on scales smaller than the Debye length, these are Landau damped, while, on scales larger than the Debye length, these are nothing but the weakly damped electron Langmuir waves and ion acoustic waves.). For a stable plasma, these eigenmodes undergo a simple exponential decay with time. On the other hand, [13] showed in a classic paper that the linear response to perturbations decays cubic exponentially due to collisions (at least at short time). Although they do not make it explicit in their paper, their analysis only applies to the continuum part of the plasma response, as we clarify in section IV. Their findings imply that the plasma echo, which arises non-linearly from the continuum response, would rapidly decay due to electron-ion collisions. The [13] prediction for the temporal decay of the continuum response is, however, much more rapid than the exponential decay of the discrete eigenmodes, identified by [9, 10]. [14] revisited the [13] result from a boundary-layer analysis of the problem, reaffirming the results of [13].
@@
It thus appears that the relaxation of weakly collisional plasmas has produced somewhat conflicting results: the super-exponential (cubic exponential) decay of the continuum response and the plasma echo according to [13] and the exponential decay of the discrete eigenmodes according to [12, 9, 10, 15, 11]. This naturally raises the following questions. Which of these results is correct? And, if both are correct, are these results simply different limits of the same response? This paper addresses these very questions. We show that the exact solution to the problem of complete perturbative (linear and second order) relaxation of a weakly collisional plasma indeed yields the above results as different asymptotic limits in time. We show that including the damping (also known as dynamical friction/drag [16]) term in the collision operator modifies the [13] prediction of cubic exponential decay of the continuum response to an exponential decay at late times and large spatial scales. We find that the [13] prediction only holds for short times, i.e., $t<1/\nu_{\rm c}$, and scales smaller than the mean free path, i.e., $\lambda<\lambda_{\rm c}=\sigma/\nu_{\rm c}$ or $\nu_{\rm c}<k\sigma$, where $\nu_{\rm c}$ is the collision frequency, $\sigma$ is the equilibrium velocity dispersion, $\lambda_{\rm c}=\sigma/\nu_{\rm c}$ is the mean free path, $k$ is the wavenumber of the response and $\lambda=2\pi/k$ is the wavelength. At long times ($t>1/\nu_{\rm c}$) and on large scales ($\lambda>\lambda_{\rm c}$ or $\nu_{\rm c}>k\sigma$), the response damps exponentially and much slower, at a rate similar to that predicted by [9, 10]. The damping is most pronounced for $\nu_{\rm c}\sim k\sigma$ or $\lambda\sim\lambda_{\rm c}$. The slow exponential decay outside the mean free path implies that the distribution function (DF) approaches the equilibrium Maxwellian form rapidly only within the mean free path; this is consistent with the assumption of the establishment of a local thermodynamic equilibrium (LTE) in the fluid approximation (large $\nu_{\rm c}$). Hence, we recover the fluid limit from a rigorous theory of weakly collisional plasmas.
@@ III Solving the linearized equations
where $\nu_{\rm c}$ is the collision frequency and $\sigma$ is the velocity dispersion of the system when it is in equilibrium. The Boltzmann equation with the LB operator is also known as the mean-field Kramers equation [see 23, for a detailed discussion of the initial-value problem for this equation in the context of weakly collisional plasmas and self-gravitating systems]. The fundamental assumption behind this form of the collision operator is that the collisions are sufficiently short-range and random. The LB operator conserves the particle number but not the linear momentum and kinetic energy during collisions. This shortcoming is well-known in the plasma literature [12, 9, 10, 11]. Despite this shortcoming, we expect the scaling behaviours of different aspects of plasma relaxation to be qualitatively robust (see section IV.2 for details).
@@ IV.1.1 Gaussian pulse
Conservation of particle number ensures that the pulse damps out as well as broadens due to collisions, as manifested by both $x$ and $v$ parts of the above response. The spatial broadening of the pulse occurs as $\sim\left(2\sigma^{2}/3\nu_{\rm c}^{2}\right){\left(\nu_{\rm c}t\right)}^{3}$ for $t\lesssim 1/\nu_{\rm c}$, and as $\sim\left(2\sigma^{2}/\nu_{\rm c}^{2}\right)\nu_{\rm c}t$ for $t\gtrsim 1/\nu_{\rm c}$, while the spatial damping occurs as the inverse of this. If the width of the initial pulse is small, i.e., $\sigma_{x,0}\ll\sigma/\nu_{\rm c}$, it damps and broadens in the former way for a long time, i.e., we are in the diffusive regime. The typical damping timescale in this regime is $\sim{\left(3\sigma^{2}_{x,0}/2\sigma^{2}\nu_{\rm c}\right)}^{1/3}$. On the other hand, a well-extended initial pulse with $\sigma_{x,0}\gg\sigma/\nu_{\rm c}$, falls in the damping regime, i.e., it quickly transitions to the latter phase of exponential damping and broadening, which occurs over a relatively long timescale of $\sim\left(\sigma^{2}_{x,0}/2\sigma^{2}\right)\nu_{\rm c}$. Since both timescales increase with $\sigma_{x,0}$ (for fixed $\nu_{\rm c}$ and $\sigma$), we see that small-scale perturbations collisionally damp away faster than large-scale ones. Moreover, since the former timescale decreases with $\nu_{\rm c}$ (for fixed $\sigma_{x,0}$ and $\sigma$) and the latter increases, perturbations damp slowly in both small and large $\nu_{\rm c}$ limits. The damping and broadening of the pulse due to collisions is visible in Fig. 2, which plots $f_{1}$ (solid lines) as a function of $x$ at different points of time as indicated. Note that the pulse slows down over time due to collisional damping. The three key features of Fokker-Planck collisional dynamics are the damping, broadening and deceleration of a pulse, all three of which are manifest in Fig. 2.
@@ IV.2 Langevin approach: diffusive and friction regimes
where we have used the uncorrelated nature of the stochastic force, i.e., the second of equations (32). Note that the velocity dispersion initially increases linearly with time, i.e., $\sigma^{2}(t)\sim 2{\cal D}\,t$. This is the essence of velocity diffusion or particle random walk in the velocity space, and is caused by the stochastic acceleration, $\xi(t)$. The late-time dynamics, which sets in after $t\sim 1/\nu_{\rm c}$, is governed by the frictional deceleration, $-\nu_{\rm c}{\rm d}x/{\rm d}t$. This friction/damping drives $\sigma^{2}$ towards the equilibrium value, $\lim_{t\to\infty}\sigma^{2}(t)=\sigma^{2}$. Thus, from the above equation, we infer that ${\cal D}=\nu_{\rm c}\sigma^{2}$, which is the celebrated Einstein relation. ${\cal D}$ is the velocity diffusion coefficient, which is related to the damping coefficient or collision frequency, $\nu_{\rm c}$, by the proportionality factor, $\sigma^{2}$. Therefore, friction and diffusion go hand-in-hand during collisions, and one cannot exist without the other. Consequently, neglecting the friction term in the collision operator [13, 14] provides an incomplete description of collisional dynamics, which may lead to erroneous conclusions about the long time decay of the response in weakly collisional plasmas.
@@ V Linear response: discrete modes
So far, we have studied the first part of the plasma response, the continuum, which encapsulates the collisionally damped streaming of particles. The second part of the response on the other hand consists of discrete modes [^3] For each $k$, there are countably infinite modes. that are collective coherent oscillations of the system. It can be computed by taking the inverse Laplace transform of $\tilde{G}_{1\kappa}^{(2)}(\omega)$ in equation (61). The resulting response involves a convolution of the continuum and discrete mode responses. We defer the computation of the combined response to the next section, and only discuss the behaviour of the discrete modes in this section. The discrete modes oscillate at frequencies that follow the dispersion relation,
@@ V.2 Lenard-Bernstein Branch
The collision operator in the Boltzmann equation involves the highest velocity derivative and is therefore a singular perturbation to the collisionless Boltzmann equation. This observation foreshadows our finding that the discrete mode spectrum for the weakly collisional case is qualitatively different from the collisionless case. Therefore, in addition to the Landau branch modes discussed above, the LB operator introduces an additional set of modes, which we term Lenard-Bernstein or LB modes. We indicate these using white dashed lines in Fig. 3. Unlike the Landau branch modes, the LB modes do not analytically continue to the Landau modes in the collisionless limit. Moreover, the LB modes are damped and non-oscillatory. As also found by [12], the frequencies of these modes are given by
@@
The LB modes play a crucial role in plasma relaxation in the large $\nu_{\rm c}/k\sigma$ limit. The damping timescale of the LB modes scales with $\nu_{\rm c}$ in the same way as that of the Landau branch modes in this limit. Also, in this limit, the damping time of the continuum response is the same as that of the $n=0$ LB mode. Hence, it is the LB modes that dictate the relaxation of plasmas in the limit of frequent collisions, or equivalently, on scales larger than the mean free path.
Vlasov simulations have demonstrated some of the phenomenology of wave damping described in this paper. [27] find in their simulation the survival of a non-oscillating and slowly decaying ‘entropy mode’ (also discussed by [21]) in the long-time limit. This bears qualitative resemblance to the LB modes [12] discussed in this paper. The long term, slow decay of the response in their simulation is very similar to that of the slowly decaying continuum and the $n=0$ LB mode that we predict (although non-linear Landau damping effects become important after approximately the bounce time or libration time for the wave-particle resonances). The pronounced damping of the response for $\nu_{\rm c}\sim k\sigma$, as predicted by our analysis, is seen by [28] in their Vlasov simulations, who note that the collision frequency suitable for preventing the “numerical recurrence of initial conditions” (the echo; see section VII for a detailed discussion of how echoes damp) is a function of the wavenumber of the perturbation. We show that this wavenumber is nothing but $\sim\nu_{\rm c}/\sigma$, the inverse of the mean free path. Perturbations tend to damp away cubic exponentially within a mean free path, and undergo slow, exponential damping beyond.